# All Patches Matter, More Patches Better: Enhance AI-Generated Image Detection via Panoptic Patch Learning

### v1

@@ 1 Introduction
The rapid evolution of generative AI models has precipitated an exponential growth of AI-generated images (AIGIs) in digital ecosystems [5, 11, 12, 10, 24, 42, 23, 38]. This proliferation raises critical concerns regarding information security and content authenticity, highlighting the critical need for AIGI detection methods to distinguish synthetic images from authentic ones. Unlike conventional classification tasks, AIGI detection operates as a ‘cat-and-mouse game’, presenting unique challenges due to: (1) continuous emergence of new generative architectures, and (2) frequent updates to existing models. Consequently, exhaustive training on all synthetic data becomes impractical [19], thus necessitating the detectors with strong generalizability.
@@
Although AIGI detection poses additional challenges for models in capturing generalizable features compared to traditional binary classification tasks, AIGIs offer a unique characteristic absent in conventional classification tasks that can be leveraged: Universal Artifact Distribution. In the context of AIGIs, discriminative features are not confined to regions with label objects; instead, synthetic images present artifacts uniformly across all patches due to the uniform production of generative models.††This work adheres to the mainstream AIGI detection setting [1, 19, 27, 15, 8, 46, 29] where the entire image is generated by AI models. This finding indicates that every image patch contains synthetic traces, establishing our principle for AIGI detection: All Patches Matter. This principle is also validated by two lines of evidence: (1) Visual analytics [29, 3] confirm pixel-level discriminative patterns in localized regions, revealing artifact presence at patch granularity; (2) recent patch-wise detectors [2, 45] show comparable performance to full-image approaches, proving individual patches’ discriminative capability. Meanwhile, while artifact variations occur across different patches, detectors capable of capturing diverse synthetic artifacts across distributed patches reduce the over-reliance on specific patches. This universal artifact capturing enhances cross-generator generalizability by mitigating detectors’ blind spots through distributed artifact aggregation. This leads to our second principle: More Patches Better.
@@
However, our counterfactual analysis of existing detectors [19, 15, 27, 8, 1] reveals an unfavorable tendency: Few-Patch Bias, involving two empirical observations and a quantitative analysis. The empirical observations: (1) detectors’ attention maps disproportionately focus on limited image patches, neglecting broader artifact patterns; (2) detectors exhibit severe patch-specific fragility, where masking merely a patch could lead to accuracy degradation by 18.7% 4.1% on average. Furthermore, by employing the causal inference tool Total Direct Effect (TDE) [33] to quantify each patch’s impact – calculated as the classification logit difference with and without that patch – we observe that the TDE distribution of naively-trained detectors is characterized by a few patches with high TDE values, while the majority of patches exhibit significantly smaller TDEs. This suggests that most patches remain underutilized and contribute minimally to the discriminative outcome. Moreover, when comparing TDE distributions across different detection methods, we find that methods with TDE distributions tending towards a more uniform distribution exhibit better generalizability. For instance, DRCT, with more high-TDE patches, performs significantly better than UnivFD.
@@
We attribute the Few-Patch Bias to the propensity of detectors as Lazy Learner [9, 40, 34, 43, 4, 31, 26, 39, 37]. Specifically, AIGI detectors exhibit curriculum learning behavior: once easily learned synthetic artifacts in certain patches are used to minimize training loss, the presence of these patches reduces the incentive to explore broader regions. We propose a framework called Panoptic Patch Learning (PPL) based on the principle ‘All Patches Matter, More Patches Better’. PPL involves: (1) Patch-wise contrastive learning, which aligns the features of different synthetic and real patches, ensuring consistent discriminative capability across all patches. (2) Random patch replacement, which randomly substitutes patches in the synthetic image with real counterparts, discourages over-reliance on limited patches and promotes a more uniform utilization of patches. Figure 1 illustrates the practical function and performance of Panoptic Patch Learning. Our main contributions are threefold:
@@
We formally propose the principle ‘All Patches Matter, More Patches Better’ for AIGI detection, demonstrating that broader artifact exploitation can effectively enhance detection.
@@
We conduct a detailed patch-wise analysis of AIGI detectors, utilizing the causal inference tool Total Direct Effect (TDE) to quantify each patch’s impact, revealing that Few-Patch Bias commonly exists in existing detectors.
@@
Based on the ‘All Patches Matter, More Patches Better’ principle, we propose Panoptic Patch Learning. Extensive experimental results validate the effectiveness of our approach.
@@ 2 Related Work
AIGI detection methods can be categorized into two general types: local and global detection methods.
@@
Local AIGI detectors utilize the image’s localized information, often in a patch-wise or pixel-wise manner, to differentiate AI-generated images from real ones. This approach is based on the premise that significant differences exist between real and synthetic images in low-level features. These detectors can be categorized into two groups: patch-wise and pixel-wise methods.
Patch-wise methods include: SSP [2] achieves remarkable performance by utilizing only a single simplest patch, while Patchcraft [45] processes the simplest and most complex patches separately by selecting patches with the highest and lowest entropy for detection, Zheng et al. [44] employs a patch-based CNN that leverages all patches to avoid selective patch sampling and aggregate patch features from an image. However, these patch-wise detectors are limited by their over-reliance on a small subset of patches, resulting in insufficient utilization of available patch information.
Pixel-wise methods include: NPR [30], which discriminates between real and AI-generated images by analyzing differences in neighboring pixel relationships; FreqNet [28] and SAFE [14], which leverage high-frequency information to detect forgeries by focusing on localized feature patterns; and ZED [3], which computes the coding cost of local regions using an entropy-based encoder and identifies AIGIs by detecting gaps in coding costs. However, these pixel-wise detectors are highly sensitive to minor variations in localized pixel relationships, which can limit their robustness in practical applications.
@@
Global AIGI detection methods leverage global information from entire images to distinguish AIGIs from real ones. CNNSpot [35] simply employs a CNN to detect AIGIs, exhibiting strong performance on seen AIGIs but suffering from poor cross-generator generalizability. UnivFD [19] addresses this limitation by utilizing a CLIP visual encoder as an AIGI feature extractor, significantly improving generalization, FatFormer [15] further improves the CLIP vision encoder’s adaptability by integrating a frequency adapter, C2P-CLIP [27] introduces a novel approach by fine-tuning CLIP with elaborately designed image-text pairs to embed the concepts of ‘real’ and ‘fake’ into the model. DRCT [1] uses a contrastive loss with hard cases to improve UnivFD’s performance, Despite the aforementioned advantages, global information does not encompass detailed artifacts of AIGIs, which severely limits their performance.
@@ 3.1 All Patches Matter, More Patches Better
The principle of All Patches Matter is built upon three key findings. (1) Theory: As every patch in synthetic images is inherently synthetic, each should contain synthetic artifacts. A series of localized region-based detection methods [2, 45] confirm that trace cues in localized patches can be used for real-synthetic discrimination, achieving substantial performance. This establishes the foundation for the principle that all patches matter, as every patch contains discriminative artifacts. (2) Visualization: Figure 4 visually illustrates the different artifact patterns present in each patch, showing that every synthetic patch has distinguishing features that differentiate it from real images. Moreover, the trace cues in different patches vary visually, confirming the diversity of artifacts. (3) Experiments: We evaluated the presence of these artifacts by inputting a randomly selected patch into the detectors, achieving an accuracy of 90% on the SDv1.4 subset of the GenImage dataset. Specifically, we replicated a single patch multiple times to form an image, ensuring that it contains only the features of that patch. The results indicate that detectors can effectively distinguish real from synthetic images even with a single random patch. The results indicate that detectors can effectively distinguish between real and synthetic images even using a single random patch. From the visualization and experiments, we recognize that artifacts in each patch are distinct, and thus, detectors are capable of capturing these artifacts, proving their ability to recognize diverse artifact patterns. This can enhance the detectors’ generalizability and robustness, leading to the principle of More Patches Better. However, our observations suggest that existing detectors do not conform to this principle.
@@ 3.2 Few-Patch Bias
Our empirical observations indicate that existing detectors often overly rely on a limited number of patches. Fig. 2 depicts the UnivFD attention map, demonstrating that the attention weight concentrates only on a few patches. After changing the ViT backbone and applying Lora fine-tuning, we find a similar observation. To further verify our hypothesis, we mask out the patches of different sizes and observe the degradation of accuracy. Fig. 3 visualizes the performance degradation of different detectors concerning the size of the masked patch. We find that masking out a patch can lead to a significant drop in accuracy, and masking out different patches will also result in different drops in accuracy.
@@
where and represents the original input image and the image with the th patch masked. By calculating the TDE for each patch, we can assess its contribution to the classification of a synthetic image.
Fig. 5 visualizes the TDE heatmap for different detectors. From top to bottom, the cases become progressively easier for the models to detect, as indicated by the increased number of active patches and a more uniform TDE distribution. We reimplemented UnivFD [19], DRCT [1] and Breaking [44] and compare them with our approach, demonstrating that a more effective method tends to discover a greater number of patches. Overall, the visual examples highlight the existing bias towards a limited number of dominant patches with higher TDE. In the methodology section, we will address this issue from a TDE perspective and explore it further with statistical analysis.
@@ 4 Methodology
Panoptic Patch Learning enhances the effectiveness of all patches through two key components: data augmentation and learning strategy. Specifically, the data augmentation technique, referred to as Random Patch Replacement(RPR), encourages the model to capture artifacts across a broader range of patches, thereby expanding the coverage of dominant patches. Following this, our learning strategy, Patch-wise Contrastive learning (PCL), ensures that all patches, both dominant and non-dominant, are brought closer in the feature space, thereby uniformizing the impact of all patches. Through the combined effect of these two components, we achieve a more comprehensive and balanced representation derived from all patches.
Random Patch Replacement encourages ‘More Patches Better’.
Random Patch Replacement (RPR) promotes the principle of ”More Patches Better” by encouraging the model to focus on learning from a greater number of detected patches. The RPR process is applied to paired images in which each reconstructed image has a corresponding ground truth . The images are partitioned into patches, and the patch replacement function is defined as:
@@
Patch-wise Contrastive Learning emphasizes ‘All Patches Matter’.
Patch-wise Contrastive Learning (PCL) operationalizes the principle of ‘All Patches Matter’ by aligning the embedding vectors of different patches, bringing patches with identical labels closer together while distancing those with different labels. We employ contrastive learning to cluster synthetic patches more closely within each batch while maintaining a margin to separate synthetic and real patches. This approach ensures that if an image contains any dominant patch with easily learnable artifacts, the model enhances its performance on the remaining patches, thus leveraging the significance of all patches. Specifically, for each batch, we utilize a margin-based contrastive loss [6] that:
where and represent the index of patch tokens within a batch. measures the Euclidean distance between the embedded patch tokens. indicates whether two patches in the pair share identical labels, thus pulling positive patch pairs (with identical patch labels) closer and pushing negative patch pairs (with different patch labels) further. The overall learning objective is a weighted combination of the cross-entropy loss and the patch-wise contrastive loss:
Analysis and Comparison.
We analyze the superiority of our approach through a quantitative analysis of our model’s performance both during and after training.
Fig. 7 (a) illustrates the loss reduction of CLIP when naively fine-tuned using LoRA compared to our proposed strategy. Both models were trained with the same batch size and learning rate on GenImage with SDv1.4 reconstructed data. The significant decrease in the loss function during the first 200 to 400 steps of naive training suggests that it is shortcutting to certain artifacts existing in the training set, limiting its generalization performance. In contrast, our method is less prone to overfitting specific artifacts.
Fig. 7 (a) illustrates the TDE distribution of UnivFD and our method. For better statistical analysis, we normalize the TDE values to a range using the exponential function . This normalization facilitates the measurement of differences between less dominant patches and the most dominant patches in the images. The figure demonstrates that a greater number of patches from our method are closer to the most dominant patches.
@@
5 Experiments
Settings.
Implementational details.
We utilized two pre-trained ViT models, CLIP [22] and DINOv2 [20], as the backbones for PPL and fine-tuned them using LoRA. During training, the input images are randomly cropped into a size of 224 224. Unless otherwise specified, in PPL, the patch size is set to 14 14, consistent with the patch size of ViT. For the Random Patch Replacement (RPR) module, each image has a probability of 0.9 for performing RPR. RPR randomly replaces of synthetic patches with real counterparts. For PCL, the weight of the contrastive loss is set to with a margin of . To achieve better performance, we add reconstructed images to the training set, following the approach of DRCT [1].
5.1 Comparison with other methods
Comparison on GenImage (Setting-I).
Tab. 1 compares PPL to other methods on GenImage. We re-implement Zheng et al. [44] and obtain the result of SAFE from [14], and the rest of the data can be sourced from C2P-CLIP[27]. We observe the following: (1) PPL consistently outperforms other methods in accuracy across various backbones. (2) The standard deviation (std) of PPL’s accuracy is significantly lower than that of the other methods, indicating greater stability across different generation methods.

### v2

@@ 1 Introduction
The rapid evolution of generative AI models has precipitated an exponential growth of AI-generated images (AIGIs) in digital ecosystems [5, 11, 12, 10, 25, 43, 24, 39, 38]. This proliferation raises critical concerns regarding information security and content authenticity, highlighting the critical need for AIGI detection methods to distinguish synthetic images from authentic ones. Unlike conventional classification tasks, AIGI detection operates as a "cat-and-mouse game", presenting unique challenges due to: (1) continuous emergence of new generative architectures, and (2) frequent updates to existing models. Consequently, exhaustive training on all synthetic data becomes impractical [20], thus necessitating detectors with strong generalizability.
@@
Although AIGI detection poses additional challenges for models in capturing generalizable features compared to traditional binary classification tasks, AIGIs offer a unique characteristic absent in conventional classification tasks that can be leveraged: Universal Artifact Distribution. In the context of AIGIs, discriminative features are not confined to regions with labeled objects; instead, synthetic images present artifacts uniformly across all patches due to the uniform production of generative models.111This work adheres to the mainstream AIGI detection setting [1, 20, 28, 16, 8, 47, 30] where the entire image is generated by AI models. This finding indicates that every image patch contains synthetic traces, establishing our first principle for AIGI detection: All Patches Matter. This principle is also validated by two lines of evidence: (1) Visual analytics [30, 3] confirm pixel-level discriminative patterns in localized regions, revealing artifact presence at patch granularity; (2) recent patch-wise detectors [2, 46] show comparable performance to full-image approaches, proving individual patches’ discriminative capability. Meanwhile, while artifact variations occur across different patches, detectors capable of capturing diverse synthetic artifacts across distributed patches reduce the over-reliance on specific patches. This universal artifact capturing enhances cross-generator generalizability by mitigating detectors’ blind spots through distributed artifact aggregation. This leads to our second principle: More Patches Better.
@@
However, our counterfactual analysis of existing detectors [20, 16, 28, 8, 1] reveals an unfavorable tendency: Few-Patch Bias, involving two empirical observations and a quantitative analysis. The empirical observations: (1) detectors’ attention maps disproportionately focus on very limited image patches, neglecting broader artifact patterns; (2) detectors exhibit severe patch-specific fragility, where masking merely a patch could lead to accuracy degradation by 18.7% 4.1% on average. Furthermore, by employing the causal inference tool TDE [33] to quantify each patch’s impact – calculated as the classification logit difference with and without that patch – we observe that the TDE distribution of naively-trained detectors is characterized by a few patches with high TDE values, while the majority of patches exhibit significantly smaller TDEs. This suggests that most patches remain underutilized and contribute minimally to the discriminative outcome, although they also contain generative artifacts. Moreover, when comparing TDE distributions across different detection methods, we find that methods with TDE distributions tending towards a more uniform distribution exhibit better generalizability. For instance, DRCT, with more high-TDE patches, performs significantly better than UnivFD. We attribute the Few-Patch Bias to the propensity of detectors as Lazy Learner [9, 41, 34, 44, 4, 31, 27, 40, 37]. Specifically, AIGI detectors exhibit curriculum learning behavior: once easily learned synthetic artifacts in certain patches are used to minimize training loss, the presence of these patches reduces the incentive to explore broader regions.
@@
To address this challenge, we argue that the key lies in following the principle: "all patches matter, more patches better", which ensures that the model does not overly rely on and shortcut to very few regions for discrimination but instead learns comprehensive and robust features across the entire image. To implement this idea, we propose a practical and effective instance called Panoptic Patch Learning (PPL), which involves: (1) Patch-wise contrastive learning, which aligns the features of different synthetic and real patches, ensuring consistent discriminative capability across all patches. (2) Random patch replacement, which randomly substitutes patches in the synthetic image with real counterparts, discourages over-reliance on limited patches and promotes a more uniform utilization of patches. Fig. 1 illustrates the practical function and performance of Panoptic Patch Learning. Our main contributions are threefold:
@@
We formally propose the principle "All Patches Matter, More Patches Better" for AIGI detection, demonstrating that broader artifact exploitation can effectively enhance detection.
@@
We conduct a detailed patch-wise analysis of AIGI detectors, utilizing the causal inference tool TDE to quantify each patch’s impact, revealing that Few-Patch Bias commonly exists in existing detectors.
@@
Based on the "All Patches Matter, More Patches Better" principle, we propose Panoptic Patch Learning. Extensive experimental results validate the effectiveness of our approach.
@@ 2 Related Work
Existing AIGI detection methods can be generally categorized into two types: local and global detection methods [28]. We will briefly introduce both of them separately below.
@@
Local AIGI detectors utilize the localized information in the image to differentiate AI-generated images from real ones. This approach is based on the assumption that significant differences exist between real and synthetic images in low-level features. These detectors can be categorized into two groups: patch-wise and pixel-wise methods.
Patch-wise methods include: SSP [2] achieves remarkable performance by using only one single simplest patch, while Patchcraft [46] processes the simplest and most complex patches separately by selecting patches with the highest and lowest entropy for detection. [45] employs a patch-based CNN that leverages all patches to avoid selective patch sampling and aggregate patch features from an image. TextureCrop [13] utilizes a sliding window to partition the input image and then ranks and selects the regions of crops that are rich in high-frequency texture information. However, these patch-wise detectors are limited by their over-reliance on a small subset of patches, resulting in insufficient utilization of available patch information.
Pixel-wise methods include: NPR [30], which discriminates between real and AI-generated images by analyzing differences in neighboring pixel relationships; FreqNet [29] and SAFE [15], which use high-frequency information to detect forgeries by focusing on localized feature patterns; and ZED [3], which computes the coding cost of local regions using an entropy-based encoder and identifies AIGIs by detecting gaps in coding costs. However, these pixel-wise detectors are highly sensitive to minor variations in localized pixel relationships, which can limit their robustness in practical applications.
@@
Global AIGI detection methods leverage information from the entire image to distinguish AI-generated images from real ones. By analyzing holistic image characteristics, these methods aim to capture global inconsistencies that may not be evident at the local level. CNNSpot [35] simply employs a CNN to detect AIGIs, exhibiting strong performance on seen AIGIs but suffering from poor cross-generator generalizability. UnivFD [20] addresses this limitation by utilizing a CLIP visual encoder as an AIGI feature extractor, significantly improving generalization, FatFormer [16] further improves the CLIP vision encoder’s adaptability by integrating a frequency adapter, C2P-CLIP [28] introduces a novel approach by fine-tuning CLIP with elaborately designed image-text pairs to embed the concepts of "real" and "fake" into the model. DRCT [1] uses a contrastive loss with hard cases to improve UnivFD’s performance, Despite the aforementioned advantages, global information does not encompass detailed artifacts of AIGIs, which severely limits their performance.
@@ 3.1 All Patches Matter, More Patches Better
The principle of All Patches Matter is supported by three key findings. (1) Theory: Since every patch in synthetic images is inherently synthetic, each patch contains synthetic artifacts. Several localized region-based detection methods [2, 46] demonstrate that trace cues within localized patches can effectively discriminate between real and synthetic images, achieving substantial performance. This evidence underpins the principle that all patches matter, as each patch carries discriminative artifacts. (2) Visualization: Fig. 2 clearly illustrates the distinct artifact patterns present in each patch, demonstrating that every synthetic patch possesses unique and identifiable features that effectively differentiate it from real images. Furthermore, the visual variation in trace cues across different patches highlights and confirms the considerable diversity of artifacts observed. (3) Experiments: We evaluated the presence of these artifacts by inputting a randomly selected patch into the detectors, achieving an accuracy of 90% on the SDv1.4 subset of the GenImage dataset. Specifically, we replicated a single patch multiple times to form an image, ensuring that it contains only the features of that patch. The results indicate that detectors can effectively distinguish real from synthetic images even with a single random patch. The results indicate that detectors can effectively distinguish between real and synthetic images even using a single random patch. From visualization and experiments, we recognize that artifacts in each patch are distinct and thus detectors are capable of capturing these artifacts, proving their ability to recognize diverse artifact patterns. This can enhance the generalizability and robustness of the detectors, leading to the principle of More Patches Better. However, our observations suggest that existing detectors do not conform to this principle.
@@ 3.2 Few-Patch Bias
Our empirical observations indicate that existing detectors often overly rely on a limited number of patches. Fig. 3(a) depicts the attention map of naively fine-tuned ViTs, demonstrating that the attention weight concentrates only on a few patches. By changing the ViT backbone and applying Lora fine-tuning, we find a similar observation. To further verify our hypothesis, we mask out the patches of different sizes and observe the degradation of accuracy. Fig. 3(b) visualizes the performance degradation of UnivFD detector concerning the size of the masked patch. We find that masking out a patch can lead to a significant drop in accuracy, and masking out different patches will also result in different drops in accuracy.
@@
where and represents the original input image and the image with the th patch masked, which we implement by setting masked patch to zero. By calculating the TDE for each patch, we can assess its contribution to the classification of a synthetic image.
Fig. 4 visualizes the TDE heatmap for different detectors. From left to right, the cases become progressively easier for the models to detect, as indicated by the increased number of active patches and a more uniform TDE distribution. We reimplemented UnivFD [20], DRCT [1] and Breaking [45] and compare them with our approach, demonstrating that a more effective method tends to discover a greater number of patches. Overall, the visual examples highlight the existing bias towards a limited number of dominant patches with higher TDE. In the methodology section, we will address this issue from a TDE perspective and explore it further with statistical analysis.
@@ 4 Methodology
Panoptic Patch Learning fulfills the idea of "All Patches Matter" and "More Patches Better" by two key components shown in Fig. 5. Specifically, the data augmentation technique, referred to as Random Patch Replacement (RPR), encourages the model to capture artifacts across a broader range of patches, thereby expanding the coverage of more patches. Following this, our learning strategy, Patch-wise Contrastive learning (PCL), ensures that all patches, both dominant and non-dominant, are brought closer in the feature space, thereby uniformizing the impact of all patches. Through the combined effect of these two components, we achieve a more comprehensive and balanced representation derived from all patches.
Random Patch Replacement encourages "More Patches Better".
Random Patch Replacement (RPR) promotes the principle of "More Patches Better" by encouraging the model to focus on learning from a greater number of detected patches. The RPR process is applied to paired images in which each reconstructed image has a corresponding ground truth . The images are partitioned into patches, and the patch replacement function is defined as:
@@
Patch-wise Contrastive Learning emphasizes "All Patches Matter".
Patch-wise Contrastive Learning (PCL) operationalizes the principle of "All Patches Matter" by aligning the embedding vectors of different patches, bringing patches with identical labels closer together while distancing those with different labels. We employ contrastive learning to cluster synthetic patches more closely within each batch while maintaining a margin to separate synthetic and real patches. This approach ensures that if an image contains any dominant patch with easily learnable artifacts, the model enhances its performance on the remaining patches, thus leveraging the significance of all patches. Specifically, for each batch, we utilize a margin-based contrastive loss [6] that:
where and represent the index of patch tokens within a batch. measures the Euclidean distance between the embedded patch tokens. defines a minimum distance threshold between negative sample pairs, thereby enhancing the model’s ability to distinguish between similar and dissimilar pairs. indicates whether two patches in the pair share identical labels, thus pulling positive patch pairs (with identical patch labels) closer and pushing negative patch pairs (with different patch labels) further. The overall learning objective is a weighted combination of the cross-entropy loss and the patch-wise contrastive loss:
@@
| Training Dataset | CNNSpot | FreDect | Fusing | UnivFD | DIRE | PatchCraft | NPR | AIDE | Ours/CLIP | Ours/DINOv2 |
| SD v1.4 | 60.11 | 56.86 | 57.07 | 55.62 | 59.71 | 56.32 | 58.13 | 62.60 | 63.94 | 66.63 |
| 8.86/98.63 | 1.37/98.57 | 0.00/99.96 | 17.65/93.50 | 11.86/95.67 | 3.07/96.35 | 2.43/100.00 | 20.33/94.38 | 17.27/99.01 | 64.65/68.12 | |
| All GenImage | 60.89 | 57.22 | 57.09 | 60.42 | 57.83 | 55.70 | 57.81 | 65.77 | 69.33 | 72.07 |
| 9.86/99.25 | 0.89/99.55 | 0.02/99.98 | 85.52/41.56 | 2.09/99.73 | 1.39/96.52 | 1.68/100.00 | 26.80/95.06 | 38.93/92.18 | 49.68/88.99 |
5 Experiments
Settings.
Implementation details.
We utilized two pre-trained ViT models, CLIP [23] and DINOv2 [21], as the backbones for PPL and fine-tuned them using LoRA. For training, the input images are randomly cropped into a size of 224 224. For testing, images exceeding 224×224 are center-cropped to this size. To achieve better performance and create paired real and fake data, we add reconstructed images with diffusion process to the training set, following the approach of DRCT [1]. To fine-tune the LORA model of CLIP or DINOv2 large with a patch-wise loss function requires 30 GB of GPU memory. Due to space limitations, additional implementation details are provided in the Appendix.
5.1 Comparison with other methods
Comparison on GenImage (Setting-I).
Tab. 1 compares PPL to other methods on GenImage. We re-implement [45] and obtain the result of SAFE from [15], and the rest of the data can be sourced from C2P-CLIP[28]. We observe the following: (1) PPL consistently outperforms other methods in accuracy across various backbones. (2) The standard deviation (std) of PPL’s accuracy is significantly lower than that of the other methods, indicating greater stability across different generation methods.